# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ 5 Numerical examples
We now use numerical continuation and bifurcation methods to corroborate the local results from §4, and to extend them to a more global picture. This includes the study of secondary bifurcations with interesting stable shapes far away from the cylinder. The numerical setup follows the stability analysis in subsection 3.3 and §4.2. Namely, we throughout fix the period $L$, and the area ${\cal A}={\cal A}(X)$, and initially use $\lambda_{2}$ as the primary bifurcation parameter and $\lambda_{1}$ and ${\cal V}(X)$ as secondary (dependent) parameters. The linearized stability of the solutions, however, is then computed for fixed ${\cal A}(X)$ and ${\cal V}(X)$ while $\lambda_{1}$ and $\lambda_{2}$ are free parameters. Due to the radius-length scaling invariance of the Helfrich energy (3) and the area and volume constraint in the dynamical problem, it is useful to introduce the reduced volume
@@
where ${\cal V}_{0}$ is the volume of the straight cylinder with the same ${\cal A}(X)$, compare [JSR93], and we normalize the energy
@@
from Fig.2, and mostly we use $L=10$. For (a) and (b) this is no significant restriction as in these cases the primary bifurcations (to pearling, buckling and coiling in (a), and to pearling and wrinkling in (b)) are of long wave type for any (sufficiently large) $L$, i.e., a different $L$ gives the same types of primary bifurcations with different axial (maximal) periods $L$ (and the wrinkling is independent of $L$). However, for (c) and sufficiently large $L$ (preferably an integer multiple of the critical period $\ell\approx 7.5$) we have a finite wave number pearling instability. Thus, besides the case $L{=}10$ (for comparison with (a) and (b)), for (c) we shall consider different multiples of 7.5 for $L$, namely $L{=}15,L{=}30$, and $L{=}45$, allowing 2, 4, and 6, pearls on the primary branch. The influence of $L$ will then be the behavior of this branch away from bifurcation, which loses stability in a bifurcation to a single pearl branch, and this BP moves closer to the primary BP with increasing $L$. This is further discussed in §LABEL:dsec.
@@
[…]
Some cylindrical solutions have similarities with axisymmetric closed vesicle solutions of (LABEL:helf0). We recall from [JSR93] or [MU24], that from the sphere the first bifurcation is to prolates (rods) and oblates (wafers). For mildly negative or positive $c_{0}$ the prolates are stable near bifurcation and the oblates unstable, and continuing the prolates they become dumbbells (two balls connected by a thin neck), from which pea shaped vesicles bifurcate which become stable as one of the ball shrinks. For sufficiently negative $c_{0}$, the oblates are stable near bifurcation, and continuing the branch they become discocytes of biconcave red blood cell shape. Here we see some reminiscent behavior, e.g., secondary bifurcations of stable wrinkles with “embedded biconcave shapes” for sufficiently negative $c_{0}$, see, e.g., Fig.7E, and secondary bifurcations to “unequal pearls” from the pearling at positive $c_{0}$ (see Figures 1012). $\rfloor$
@@
[…]
@@
    Naturally, we want $E$ to decrease monotonously in $t$, cf.(LABEL:helf1b), which if violated gives an easy indicator of poor numerics. The monotonous decrease of $E$ often requires rather small stepsizes and tolerances, and a result is that in many examples a single numerical flow to steady state is as expensive as the computation of the full BD of steady states, which we take as a strong argument for our steady state continuation and bifurcation approach.
@@ 5.1 $c_{0}=0$
For $L=10$ and $c_{0}=0$, Corollary 3.2 gives that the cylinder destabilizes at $\lambda_{2}\approx-0.95$ to pearling, $(m,n)=(1,0)$, and at $\lambda_{2}\approx 1.19$ to modes $(m,n)=(1,1)$. Figure 4 shows the BDs and samples starting in (a1) with the energy $E$ over the reduced volume $v$.
[…]
@@
Figure 5: $c_{0}=0$, comparison of numerics and AE predictions (lines with *), (a) pearling, (b) coils (orange) and buckling (green). Left and center: Amplitude over reduced volume $v$, and over $\lambda_{2}$. Right: $\lambda_{1}$ over $\lambda_{2})$.
[…]
@@
[…]
@@
Figure 6: Numerical flows from perturbations of states from Fig.4; initial and ’final’ state, and evolution of $E$, $\lambda_{1,2}$ (only $\lambda_{2}$ in (a) and (c2)), and deviations from intial area and volume; very slow evolution in (c2).
The first two flows start near stable steady states $(X_{0},\lambda_{1,0},\lambda_{2,0})$, and as expected flow back to near $X_{0}$ rather quickly, with partly significant changes of $(\lambda_{1},\lambda_{2})$. However, the third flow (c1) is somewhat surprising. State C from Fig.4 is linearly unstable, but the perturbed $X_{0}$ flows to another buckling by adjusting $\lambda_{1,2}$ at the increased area and volume, where this buckling is stable. On the other hand, in (c2) for the (linearly) area and volume preserving initial perturbation into the most unstable direction (which point towards coils), the behavior is different: The solution now converges to a coil, with a rather small flow throughout; $\lambda_{1}$ hardly changes at all, $\lambda_{2}$ only very slightly, and the convergence is very slow (since the leading eigenvalue of the target coil is negative but very close to zero). For all flows, we can use the final state as an initial guess for steady state continuation (e.g., in $\lambda_{2}$), and the obtained numerical steady states then are (volume preserving) linearly stable.
[…]
@@
[…]
@@
Refer to caption Refer to caption Refer to caption
Figure 7: Basic BD (a) for $c_{0}=-1$, $L=10$, and samples (b). (a1) Energy $E$ over $v$; (a2) $\lambda_{1}$ over $\lambda_{2}$ for the wrinkling D and ’embedded oblates’ E. (c) Comparison with AE prediction for primary wrinkling.
5.2 $c_{0}=-1$
For (strongly) negative $c_{0}$, the wrinkling branches and their secondary bifurcations become relevant. Fig.7(a,b) shows basic results for $c_{0}{=}-1$ and $L{=}10$, and (c) a comparison with the AE predictions from (62) for wrinkling, again with very good agreement, cf. Remark 5.1. As before, at negative $\lambda_{2}\approx-2.254$ and $\lambda_{1}\approx 1.5$ (cf.Fig.2), $\mathcal{C}_{r}$ looses stability to pearling, but now at positive $\lambda_{2}=3/2$ and $\lambda_{1}=-9/4$ it looses stability to the wrinkling branch (samples D1,2), while the bifurcation to buckling and coiling is at a later $(\lambda_{1},\lambda_{2})\approx(-3.77,3.021)$ and yields initially unstable branches. The wrinkling is initially stable, but at $\lambda_{2}\approx 1.6$ loose stability to “oblates embedded in wrinkling”, see sample E, and Remark 5.
Figure 8 shows some numerical flows from states from Fig.7, focusing on wrinkling perturbed according to (67) (pearling is again stable, also under rather large perturbations). D2 at its given $({\cal A}_{0},{\cal V}_{0})$ is linearly (weakly) unstable, and the perturbation yields a slow evolution to an oblate wrinkling, with a very small change in $E$. In (b) and (c) we perturb a coiling and a buckling, and in both cases flow to wrinkling, and the bottom line of Fig.8 and similar experiments is that (stable) wrinkling and oblate wrinkling have rather large domains of attraction.
@@ 5.2 $c_{0}=-1$
Lastly, we consider positive $c_{0}$, specifically $c_{0}=0.48$. By Corollary 3.2, the cylinder destabilizes in a finite wave length pearling instability, and in a long wave instability to coiling and buckling. The finite pearling period is $\ell:=\ell_{p}\approx 7.5$, and thus we will be mostly interested in length $L$ as multiples of $7.5$.
However, for comparison with the cases $c_{0}=0$ and $c_{0}=-1$ we first again choose $L=10$, and in Fig.9 give the BD of buckling and coils. The coefficients for the AE description (64) now are $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$, and in agreement with Fig.3 the buckling are now stable at bifurcation and the coils unstable. Moreover, we find interesting 2ndary bifurcations on both branches, where the buckling looses (coiling gains) stability, and these two BPs are connected by a stable intermediate branch, see sample C. The comparison with the AEs in (c) again shows the agreement close to bifurcation, but naturally the (cubic order) AEs cannot capture the fold of the coiling in $\lambda_{2}$.
@@
[…]
@@
Figure 10 shows the BD and samples for $L=15$. To the right of $v=1$ bifurcates the $m=2$ pearling branch (at $\lambda_{2}\approx-0.243$), and is initially stable (sample A1). However, shortly after ($\lambda_{2}\approx-0.3756$) there is a secondary bifurcation to a stable branch with double the period, sample B. On the other hand, the primary $m=1$ pearling branch, sample C, (bifurcating at $\lambda_{2}\approx-0.421$) is unstable throughout. The $(m,n)=(1,1)$ coiling (sample D) and buckling (sample E) bifurcate at $\lambda_{2}\approx 0.138$, and again the coiling is stable, and stays stable to rather small $v$, while the buckling is unstable. Wrinkling does not occur in the parameter regime considered; however, the (compared to $L=10$ before) slightly larger $L=15$ lets the pearling and coiling/buckling BPs move closer together, which is why we chose to present the $(m,n)=(2,1)$ coiling F (unstable). Figure 11 shows the flow from a perturbation of A2 to a B type solution but rather close to the BP. Therefore, the convergence to steady state is again very slow, as near bifurcation the stable B branch has a leading eigenvalue close to 00.
[…]
@@
The main point of Fig.10 is that the choice of $c_{0}=0.48$ together with $L=15$ yields the destabilization of the cylinder to the $m=2$ pearling, but these solutions are only stable in a somewhat narrow range (in $v$, say). To further assess this, in Fig.12 we consider the primary pearling with $\ell=7.5$ on twice the domain $L=30$. The result shows that the range of stable $\ell=7.5$ pearls shrinks with increasing $L$, and this continues for further increase of $L$, yielding the following dependence of the reduced volume $v_{s}$ of single pearling BPs from primary $\ell=7.5$ pearling on $L$:

### v2

@@ 5 Numerical examples
We now use numerical continuation and bifurcation methods to corroborate the local results from §4, and to extend them to a more global picture. This includes some secondary bifurcations with interesting stable shapes away from the straight cylinder. The numerical setup follows the stability analysis in subsection 3.3 and §4.2. Namely, we fix a period $L$ (e.g., $L=10$, $L=15$ or $L=30$, see also Remark 3.3), and the area ${\cal A}={\cal A}(X)$, and initially use $\lambda_{2}$ as the primary bifurcation parameter and $\lambda_{1}$ and ${\cal V}(X)$ as secondary (dependent) parameters. The linearized stability of the solutions, however, is then computed for fixed ${\cal A}(X)$ and ${\cal V}(X)$ while $\lambda_{1}$ and $\lambda_{2}$ are free parameters. Due to the radius-length scaling invariance of the Helfrich energy (1) and the area and volume constraint in the dynamical problem, it is useful to introduce the reduced volume
@@
where ${\cal V}_{0}$ is the volume of the straight cylinder with the same ${\cal A}(X)$, compare [JSL93], and we normalize the energy
@@
from Fig.2, and mostly we use $L=10$. For (a) and (b) this is no significant restriction as in these cases the primary bifurcations (to pearling, buckling and coiling in (a), and to pearling and wrinkling in (b)) are of long wave type for any (sufficiently large) $L$, i.e., a different $L$ gives the same types of primary bifurcations with different axial (maximal) periods $L$ (and the wrinkling is independent of $L$). However, for (c) and sufficiently large $L$ (preferably an integer multiple of the critical period $\ell\approx 7.5$) we have a finite wave number pearling instability. Thus, besides the case L=10L{=}10 (for comparison with (a) and (b)), for (c) we shall consider different multiples of 7.5 for $L$, namely L=15,L=30L{=}15,L{=}30, and L=45L{=}45, allowing 2, 4, and 6, pearls on the primary branch. The influence of $L$ will then be the behavior of this branch away from bifurcation, which loses stability in a bifurcation to a single pearl branch, and this BP moves closer to the primary BP with increasing $L$. This is further discussed in §6. Additionally, in §5.4 we give an outlook on case (d) $c_{0}=0.75$, from Fig.2, where ${\cal C}_{L}$ is unstable for all $\Lambda$.
@@
Remark 5.1.
Some cylindrical solutions have similarities with axisymmetric closed vesicle solutions of (4). We recall from [JSL93] or [MU24b], that from the sphere the first bifurcation is to prolates (rods) and oblates (wafers). For mildly negative or positive $c_{0}$ the prolates are stable near bifurcation and the oblates unstable, and continuing the prolates they become dumbbells (two balls connected by a neck), from which pea shaped vesicles bifurcate which become stable as one of the ball shrinks. For sufficiently negative $c_{0}$, the oblates are stable near bifurcation, and continuing the branch they become discocytes of biconcave red blood cell shape. Here we see some reminiscent behavior: For sufficiently negative $c_{0}$ we find secondary bifurcations of stable wrinkles with “embedded biconcave shapes”, see, e.g., Fig.7E. For sufficiently positive $c_{0}$, the primary pearling bifurcation leads to “strings of dumbbells” rather than strings of (round) pearls, see, e.g., samples C in Figures 11 and 13, or the samples A1–A3 in Fig.14. $\rfloor$
@@
Our numerical flows $t\mapsto(X(t),\Lambda(t))=\Psi_{\text{num}}(t;h)X_{0}$ are ad hoc approximations with a time stepsize $h$ and tolerance tol (see Appendix B), and we summarize the results from our flow experiments as follows:
@@
    Naturally, we want $E$ to decrease monotonously in $t$, cf.(6), which if violated gives an easy indicator of poor numerics. The monotonous decrease of $E$ often requires rather small stepsizes and tolerances, and a result is that in many examples a single numerical flow to steady state is as expensive as the computation of the full BD of steady states, which we take as a strong argument for our steady state continuation and bifurcation approach.
  4. 4.
    For $c_{0}=0.75$, where ${\cal C}_{L}$ is always unstable and hence no stable primary bifurcations occur, we can numerically flow to stable finite wavelength pearling, which we subsequently continue in parameters, see Fig.14. However, this is only an outlook, and we leave a detailed and systematic study of this for future work.
@@ 5.1 $c_{0}=0$
For $L=10$ and $c_{0}=0$, §3.3 gives that the cylinder destabilizes at $\lambda_{2}\approx-0.95$ to pearling $(k_{1},0)$, and at $\lambda_{2}\approx 1.19$ to coiling/buckling $(k_{1},1)$. Figure 4 shows the BDs and samples starting in (a1) with the energy $E$ over the reduced volume $v$.
(a1) (a2)
@@
Figure 5: $c_{0}=0$, comparison of numerics and AE predictions (lines with *), (a) pearling, (b) coils (orange) and buckling (green). Left and center: Amplitude over reduced volume $v$, and over $\lambda_{2}$. Right: $\lambda_{1}$ over $\lambda_{2}$.
Remark 5.2.
@@
yields $\frac{{\rm d}}{{\rm d}T}A=A(0.61-2.96A^{2})$ and $\alpha_{2}=0.66A^{2}-1$, and that for the wrinkling $u=\varepsilon A\psi_{01}+{\cal O}(\varepsilon^{2})$, $\lambda_{2}=\lambda_{2}(0,1)+\varepsilon^{2}$, $\lambda_{1}=\lambda_{1}(0,1){+}\alpha_{2}\varepsilon^{2}{+}{\cal O}(\varepsilon^{4})$ yields the universal AE (65), i.e., $\frac{{\rm d}}{{\rm d}T}A=A(3{-}15.19A^{2})$ and $\alpha_{2}=\frac{27}{2}A^{2}{-}1$. For both, pearling and wrinkling, we find an excellent match with the numerics. $\rfloor$
@@
Figure 6: Numerical flows from perturbations of states from Fig.4; initial and ’final’ state, and evolution of $E$, $\lambda_{1,2}$ (only $\lambda_{2}$ in (a) and (c2)), and deviations from initial area and volume; very slow evolution in (c2).
The first two flows start near stable steady states $(X_{0},\lambda_{1,0},\lambda_{2,0})$, and as expected flow back to near $X_{0}$ rather quickly, with partly significant changes of $(\lambda_{1},\lambda_{2})$. However, the third flow (c1) is somewhat surprising. State C from Fig.4 is linearly unstable, but the perturbed $X_{0}$ flows to another buckling by adjusting $\lambda_{1,2}$ at the increased area and volume, where this buckling is stable. On the other hand, in (c2) for the (linearly) area and volume preserving initial perturbation into the most unstable direction (which point towards coiling), the behavior is different: The solution now converges to coiling, with a rather small flow throughout; $\lambda_{1}$ hardly changes at all, $\lambda_{2}$ only very slightly, and the convergence is very slow (since the leading eigenvalue of the target coiling is negative but very close to zero). For all flows, we can use the final state as an initial guess for steady state continuation (e.g., in $\lambda_{2}$), and the obtained numerical steady states then are (volume preserving) linearly stable.
Table 1: Spectral data ($n_{\text{unst}}$=number of ($V$-preserving) unstable directions, $\mu_{\text{crit}}$=least stable/most unstable eigenvalue), perturbation parameter $\delta$ ($\xi{=}1$ throughout), and results for DNS in Fig.6; $A_{0}{=}20\pi,V_{0}{=}10\pi$. Initial perturbations for A–C of type (72), and for C of type (73). $(\lambda_{1},\lambda_{2})_{0}$ are the values at the (perturbed) steady state.
@@
5.2 $c_{0}=-1$
For (strongly) negative $c_{0}$, the wrinkling branches and their secondary bifurcations become relevant. Fig.7 shows basic results for $c_{0}{=}-1$ and L=10L{=}10. As before, at negative $\lambda_{2}\approx-2.254$ and $\lambda_{1}\approx 1.5$ (cf.Fig.2), $\mathcal{C}_{r}$ loses stability to pearling, but now at positive $\lambda_{2}=3/2$ and $\lambda_{1}=-9/4$ it loses stability to the wrinkling branch (samples D1,2), while the bifurcation to buckling and coiling is at a later $(\lambda_{1},\lambda_{2})\approx(-3.77,3.021)$ and yields initially unstable branches. The wrinkling is initially stable, but at $\lambda_{2}\approx 1.6$ loses stability to “oblates embedded in wrinkling”, see sample E, and Remark 5.
@@ 5.2 $c_{0}=-1$
Figure 7: Basic BD for $c_{0}=-1$, $L=10$, and samples.
Figure 8 shows some numerical flows from states from Fig.7, focusing on wrinkling perturbed according to (72) (pearling is again stable, also under rather large perturbations). D2 at its given $({\cal A}_{0},{\cal V}_{0})$ is linearly (weakly) unstable, and the perturbation yields a slow evolution to an oblate wrinkling, with a very small change in $E$. In (b) and (c) we perturb a coiling and a buckling, and in both cases flow to wrinkling, and the bottom line of Fig.8 and similar experiments is that (stable) wrinkling and oblate wrinkling have rather large domains of attraction. For completeness, Fig.9 shows a comparison with the AE predictions from (65) for wrinkling, again with very good agreement, cf. Remark 5.1.
@@
Refer to caption Refer to caption Refer to caption
Figure 9: Comparison with AE prediction for primary wrinkling.
@@
We now consider $c_{0}\in(1/4,1/2)$, specifically $c_{0}=0.48$, cf. Fig.2(c). By Corollary 3.7, the cylinder destabilizes in a finite wavelength pearling instability, and in a long wave instability to coiling and buckling. The finite pearling period is $\ell:=\ell_{p}\approx 7.5$, and thus we will be mostly interested in length $L$ as multiples of $7.5$.
However, for comparison with the cases $c_{0}=0$ and $c_{0}=-1$ we first again choose $L=10$, and in Fig.10 give the BD of buckling and coiling. The coefficients for the AE description (67) now are $(a,b_{1},b_{2})=(0.39,-0.083,0.56)$, and in agreement with Fig.3 the buckling is now stable at bifurcation and the coiling is unstable. Moreover, we find interesting secondary bifurcations on both branches, where the buckling loses (coiling gains) stability, and these two BPs are connected by a stable intermediate branch, see sample C. The comparison with the AEs in (c) again shows the agreement close to bifurcation, but naturally the (cubic order) AEs cannot capture the fold of the coiling in $\lambda_{2}$.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(c) (d)
Refer to caption Refer to caption Refer to caption Refer to caption
@@
(a1) (a2)
@@
Figure 11 shows the BD and samples for L=15L{=}15. To the right of v=1v{=}1 bifurcates the $k_{2}{\approx}0.838$ pearling branch (at $\lambda_{2}{\approx}-0.243$), and is initially stable (sample A1). However, shortly after ($\lambda_{2}{\approx}-0.3756$) there is a secondary bifurcation to a stable branch with double the period, sample B. On the other hand, the primary m=1m{=}1 pearling branch, sample C, (bifurcating at $\lambda_{2}{\approx}-0.421$) is unstable throughout. [^6] At first sight, C may look similar to the $m=2$ pearling A1 and A2, but the difference is that C has one weak and one stronger constriction, which intuitely makes C one period in a string of dumbbells; see also C in Fig. 13, and Remark 5. The (m,n)=(1,1)(m,n){=}(1,1) coiling (sample D) and buckling (sample E) bifurcate at $\lambda_{2}{\approx}0.138$, and again the coiling is stable, and stays stable to rather small $v$, while the buckling is unstable. Wrinkling does not occur in the parameter regime considered; however, the (compared to $L=10$ before) slightly larger $L=15$ lets the pearling and coiling/buckling BPs move closer together, which is why we chose to also present the $(m,n)=(2,1)$ coiling F (unstable). Figure 12 shows the flow from a perturbation of A2 to a B type solution but rather close to the BP. Therefore, the convergence to steady state is again very slow, as near bifurcation the stable B branch has a leading eigenvalue close to 0.
@@
(a) (b)
Figure 13: $c_{0}=0.48$, $L=30$. The $m=4$ (period $\ell=7.5$) pearling A now loses stability at BP1 much closer to its primary bifurcation, to a single “localized” pearling branch, sample B. Further branches: C: $m=2$ ($\ell=15$) primary pearling; E: $m=3$ ($\ell=10)$ primary pearling; D: 2ndary pearling from 2nd BP on $m=1$ branch. F: primary $(m,n)=(1,1)$ coiling.
The main point of Fig.11 is that the choice of $c_{0}=0.48$ together with $L=15$ yields the destabilization of the cylinder to the $k_{2}\approx 0.838$ pearling, but these solutions are only stable in a somewhat narrow range (in $v$, say). To further assess this, in Fig.13 we consider the primary pearling with $\ell=7.5$ on twice the domain $L=30$. The result shows that the range of stable $\ell=7.5$ pearling shrinks with increasing $L$, and this continues for further increase of $L$, yielding the following dependence of the reduced volume $v_{s}$ of single pearling BPs from primary $\ell=7.5$ pearling on $L$: